# A hydrodynamic approach to reproduce multiple spinning vortices in horizontally rotating three-dimensional liquid helium-4

### v1

@@ 4 Discussion
A description of the dynamics of liquid helium-4 at cryogenic temperatures based on classical fluid dynamics yields Eqs.  (1) and (2), which are similar to the conventional two-fluid model derived by combining the nonlinear $\rm Schr\ddot{o}dinger$ equation for bosons, that is, the Gross-Pitaevskii equation, with the thermodynamic relation Gibbs-Duhem equation. In addition, our two-fluid model includes a conservation term for angular momentum related to the rotation of classical fluid particles. The fact that we could reproduce multiple spinning vortices, such as quantum vortices, using a method based on classical fluid mechanics would be groundbreaking because multiple spinning vortices and their lattice phenomena in liquid helium-4 were thought to originate purely from quantum mechanics. The simulation results showed that the individual vortices were thinner and that there were more vortices in the high-resolution simulation than in the low-resolution simulation. This is reasonable because if we focus on the cross section at z = 0, the number of vortices observed on the cross section should be a few hundred according to Feynman’s law in two-dimensional cases [7]. In addition, the vortex centers of quantum vortices are essentially of the order of angstroms (although quantum vortices are thought to interact even at large distances). Therefore, the results of the high-resolution simulation were closer to the theoretical solution.
@@ 5 Conclusion
In this paper, we report a 3D simulation of the rotating liquid helium-4 using a two-fluid model with spin-angular momentum conservation using SPH. The conventional two-fluid model was derived by combining the nonlinear Schrodinger equation for bosons with the thermodynamic Gibbs-Duhem equation. In short, the usual two-fluid model is obtained by making macro-corrections to the microscopic relational equations. In contrast, our model is derived from the particle approximation of an inviscid fluid with residual viscosity in part. Despite the fully classical mechanical picture, the resulting system equations were consistent with those of the conventional two-fluid model. Specifically, we consider the bulk liquid helium-4 to be an inviscid fluid, assuming that the viscous fluid component remains at finite temperatures. As the temperature decreased, the amount of the viscous fluid component decreased, ultimately becoming a fully inviscid fluid at absolute zero. Weak compressibility is assumed to express the volume change because some helium atoms do not render fluid owing to BECs or change states owing to local thermal excitation. Using explicit SPH (smoothed particle hydrodynamics), one can solve the governing equations for an incompressible fluid, simultaneously reproducing density fluctuations and describing the fluid in a many-particle system. We assume the following fluid-particle duality: a hydrodynamic interfacial tension between the inviscid and viscous components or a local interaction force between two types of fluid particles. The former can be induced in the horizontal direction when non-negligible non-uniformity of the particles occurs during forced 2D rotation, and the latter is non-negligible when the former is negligible. We performed a large-scale simulation of 3D liquid helium forced to rotate horizontally using 32 GPUs. Compared with the low-resolution calculation using 2.4 million particles, the high-resolution calculation using 19.6 million particles showed spinning vortices close to those of the theoretical solution. In the future, the introduction of quantum mechanical recombination among vortices will capture the behavior of quantum fluids. We obtained a promising venue to establish a practical simulation method for bulk liquid helium-4.

### v2

@@ 4 Discussion
A description of the dynamics of liquid helium-4 at cryogenic temperatures based on classical fluid dynamics yields Eqs. (1) and (2), which are similar to the conventional two-fluid model derived by combining the nonlinear $\rm Schr\ddot{o}dinger$ equation for bosons, that is, the Gross–Pitaevskii equation, with the thermodynamic relation Gibbs–Duhem equation. In addition, our two-fluid model includes a conservation term for angular momentum related to the rotation of classical fluid particles. The fact that we could reproduce multiple spinning vortices, such as quantum vortices, using a method based on classical fluid mechanics would be groundbreaking because multiple spinning vortices and their lattice phenomena in liquid helium-4 were thought to originate purely from quantum mechanics. The simulation results showed that the individual vortices were thinner and that there were more vortices in the high-resolution simulation than in the low-resolution simulation. This is reasonable because if we focus on the cross section at z = 0, the number of vortices observed on the cross section should be a few hundred according to Feynman’s law in two-dimensional cases [7]. In addition, the vortex centers of quantum vortices are essentially of the order of angstroms (although quantum vortices are thought to interact even at large distances). Therefore, the results of the high-resolution simulation were closer to the theoretical solution.
@@
Figure 8: Schematic explanations of the image analysis to count the number of vortices: (a) an example of the volume rendering, showing the intensity of the rotational force at time 0.32 s in the lower resolution, which corresponds to Fig. 4(c), and (b) the intensity distribution of the rotational force at $z=0$ plane, which corresponds to the scalar distribution shown in (a). (c) the correlation between the number of observed vortices and the number of particles, and (d) the relationship between the average vortex size and the number of particles, at 0.32 and 0.79 s.
Several experimental studies [17, 73] on rotated cryogenic liquid helium-4 have reported peculiar phenomena compared to those observed in rotation problems in classical hydrodynamics. Specifically, multiple quantum vortices emerge and spin around their axes parallel to the vertical center axis of the cylindrical vessel, eventually forming a regular lattice. The lattices rotate around the cylindrical axis while maintaining a constant relative position, which is called rigid-body rotation. Of particular interest is the emergence of multiple independent spinning nondissipative vortices; this phenomenon in a 3D problem was numerically reproduced for the first time in this study. To clarify the implications of the findings of this study and for future tasks, we investigated the dependence of the number and average size of vortices on the numerical resolution, that is, the number of fluid particles, at two different elapsed times. The number of vortices was measured in the following manner. Because vortices exhibit vertical structures, we measured the number of vortices that crossed the $z=0$ plane. We identified and labeled the vortices from the intensity distribution of the rotational forces exerted by the respective vortices on the $z=0$ plane using the ImageJ software [74, 75, 76]. Figures 8(a) and (b) present this process; (a) is an example snapshot of the volume rendering of the simulation, showing the intensity of the rotational force at time of 0.32 s in the lower resolution simulation with 2.4 million particles; this subfigure corresponds to Fig. 4(c); (b) shows the scalar distribution of the intensity of the rotational force in the $z=0$ plane in (a). Figures 8(a) and (b) clearly show that the edges of each vortex can be identified because the intensity of the rotational force is maximized at the edges of the vortices. Finally, Fig. 8(c) shows the correlation between the numbers of observed vortices and particles, and Fig. 8(d) shows the relationship between the average vortex size and number of particles at 0.32 and 0.79 s.
The significant observations are summarized as follows. First, vortices tend to move freely or separately, absorbing neighboring smaller vortices and changing into a smaller number of larger vortices; the total number of vortices decreases with time. In brief, local interactions among vortices are classical. In practice, fewer vortices were observed, as shown in Fig. 8(c) for the triangular symbol (at 0.79 s) compared to that for the cross symbol (at 0.32 s). However, it should be emphasized that independent vortices rarely disappear unlike in ordinary classical fluids. In summary, multiple spinning, independent, and non-dissipative vortices emerged for horizontally rotated 3D cryogenic helium-4 at a lower resolution; however, local interactions among the vortices were classical, except for their non-dissipative nature.
However, this phenomenon was observed less frequently as the resolution increased. In Fig. 8(d), as the resolution increases from low (2.4 million particles) to medium (8.62 million particles) to high (19.6 million particles), the average size of the vortices decreases and no significant difference in size is observed between two different cases of elapsed time. Specifically, the average size of the vortices can be considered to reach a steady state over time. First, considering that the proposed method does not introduce constraints on the topology changes in the vortices, the interactions between the vortices inevitably exhibit classical fluidity. Nevertheless, as the resolution increases, the vortices become smaller. This can be described as follows. In our simulations, the diameter of the fluid particles decreased as the resolution increased, but the value of the parameter $C_{\omega}$, which determines the magnitude of the angular velocity, was set to be inversely proportional to the size of the particle such that the rotational force at the edge of the individual fluid particles, expressed as their product as $r|\bm{\mathbf{\omega_{0}}}|$, remained the same. This ensures that the torque (rotational force) on each fluid particle is constant regardless of the resolution. Consequently, although the number of particles required to form a vortex remained constant, the average size of the vortices decreased because the size of the constituent particles decreased. Then, the smaller sized vortices with the same magnitude of rotational force are distributed over the entire domain. Because the average size of the vortices decreases, the frequency of vortex-vortex interactions decreases; a higher number of multiple rotating vortices with small sizes can exist compared to the low-resolution case.
As we will see later, the average size of the vortices at 19.6 million particles is comparable to that of the average vortex-vortex distance derived from Feynman’s rule in quantum hydrodynamics; from a local approximation point of view, it can be considered approximately equivalent to the velocity field generated by each quantum vortex. From a phenomenological perspective, the fact that a number of multiple independent vortices can exist with smaller sizes comparable to the velocity field of quantum vortices as the resolution increases suggests that vortices become more quantum mechanical as the resolution increases. In other words, our two-fluid model might directly describe the microscopic behavior as the resolution increases. There are two possible explanations for this observation. The first possibility is that this claim is incorrect, and that artificial control of the vortex topology is still required. In quantum mechanics, circulation is quantized. Therefore, we may need to introduce a constraint such as the Schwarz equation to maintain the circulation or each vortex constant, leading to a change in the system into a quantum fluid system. The second possibility is that such a constraint can be introduced automatically. According to the Young-Laplace equation, the surface tension of a classical fluid is proportional to its curvature. As the vortex becomes smaller, its surface tension increases, which may solidify the vortex and act as a type of topological control. To validate this, simulations with the same physical conditions and at higher resolutions must be performed.
Based on the results of this study, the vortices have more opportunities to interact with each other over longer periods. Consequently, the vortices gradually integrate and decrease to the same degree as the ideal number of vortices calculated by Feynman’s rule, which is equal to 284.7 in the corresponding 2D case (refer to Ref. [7] for specific calculations) with 19.6 million particles in Fig. 8(c). However, even in this case, the number of vortices decreased with time. Thus, the system did not reach a steady state with respect to the number of vortices in the time direction. Among the two possibilities, in the former, a comparison with the theoretical value of Feynman’s rule must be discussed in detail after the corresponding 3D physical model of circulation control is introduced, and the number of vortices becomes stable in the time direction. By contrast, in the latter, higher resolution simulations are required. In Fig. 8(c), the curve with the triangular symbol is exponential relative to the computational resolution, whereas the curve with the cross symbol is linear with respect to the computational resolution. Therefore, we can expect to reach a stationary state in the time direction when using tens to 100 million particles for the entire domain without introducing such a topology control model. In the former case, developing a 3D vortex dynamics model based on classical fluid mechanics as an alternative to Schwarz’s rule [13, 14] is another approach. Our results for the 2D problem have been presented in previous studies [7, 8], where we proposed a model that balances the Magnus force generated by each rotating vortex with the repulsive force as a vortex-vortex interaction, which prevents the vortices from moving closer than necessary. Specifically, by calculating the vortex circulation value and introducing a numerical model in which the vortex dynamics model becomes effective when the circulation exceeds a certain criterion, we successfully controlled the equal arrangement and circulation of vortices [7, 8]. Hence, in principle, it is possible to develop similar dynamic vortex models for 3D problems. However, several issues need to be resolved, such as the physical correspondence between the classical vortex dynamics model and Schwarz’s rule [13, 14], including whether they are in one-to-one correspondence. Moreover, in future research, our 2D vortex dynamics model should be extended to changes in 3D topology.
For 19.6 million particles, 500 fluid particles were arranged in each horizontal direction, as mentioned previously. Because the diameter of the disk was 0.2 cm (${\rm=2.0\times 10^{-3}m}$), the diameter of a particle is ${\rm 4\times 10^{-6}m}$ or 4 micrometers. If we assume 10 particles per vortex in each horizontal direction, the size of a single vortex would be approximately 40 micrometers. In contrast, the scale of a vortex core is on the order of angstroms. However, because the effective range of a vortex is determined by the Biot–Savart law, which describes the velocity field around the vortex, the area of influence of the vortex is considered to be its effective size. In our case, the average vortex distance was calculated using Feynman’s rule, which yielded approximate ${\rm 10^{-4}m}$ [7]. The corresponding number of vortices is shown by the dotted line in Fig. 8(c). It is reasonable to understand that the vortex observed does not correspond to an actual quantum vortex core, but to the effective velocity field created by the quantum vortex.
In summary, the next task is to control the circulation of vortices by ensuring an appropriate vortex–vortex interaction to stabilize the number of vortices in the time direction. As desctibed above, this may be achieved by introducing an artificial circulation control model by referring to Schwarz’s rule, or by simply increasing the resolution, expecting that an increase in the strength of the surface tension would solidify the vortex. We regard the latter case as promising for the following reasons. In Fig. 8(c), the curve with the triangular symbol is exponential relative to the computational resolution, whereas that with the cross symbol is linear with respect to the computational resolution. Extrapolating these results enables us to expect that the vortices may reach an approximately stationary state in the time direction when tens to 100 million particles are used in the entire domain. Another approach is to develop a 3D vortex dynamics model based on classical hydrodynamics, as an alternative to Schwarz’s rule. We achieved some success in the 2D calculations using models representing this effect in the Magnus force and vortex repulsion models [7], as described above. However, several issues need to be resolved, such as the physical correspondence between the classical vortex dynamics model and Schwarz’s rule and how to extend our 2D vortex dynamics model to changes in 3D topology. Nevertheless, this approach should be examined after checking whether the second approach is valid.
From an engineering perspective, our goal is not to propose a first-principles model, but to enable the simulation of practical bulk liquid helium; we only have to capture the dynamic interactions among vortices accurately enough to reproduce macroscopic quantum effects. In practice, a direct comparison between the internal structure within the healing length of the quantum vortex and the vortex in our method requires at least several hundred times more particles in each direction, even when using a coarse-grained approach. However, this was beyond the scope of this study. At present, this is the first study to numerically reproduce the dynamics of rotated cryogenic 3D liquid helium-4 in bulk form using a classical hydrodynamic approach. Accordingly, it is scientifically significant that we observed multiple spinning nondissipative vortices in the 3D problem of horizontally rotated cryogenic liquid helium-4.
@@ 5 Conclusion
In this paper, we report a 3D simulation of the rotating liquid helium-4 using a two-fluid model with spin-angular momentum conservation using SPH. The conventional two-fluid model was derived by combining the nonlinear Schrodinger equation for bosons with the thermodynamic Gibbs–Duhem equation. In short, the usual two-fluid model is obtained by making macro-corrections to the microscopic relational equations. In contrast, our model is derived from the particle approximation of an inviscid fluid with residual viscosity in part. Despite the fully classical mechanical picture, the resulting system equations were consistent with those of the conventional two-fluid model. Specifically, we consider the bulk liquid helium-4 to be an inviscid fluid, assuming that the viscous fluid component remains at finite temperatures. As the temperature decreased, the amount of the viscous fluid component decreased, ultimately becoming a fully inviscid fluid at absolute zero. Weak compressibility is assumed to express the volume change because some helium atoms do not render fluid owing to BECs or change states owing to local thermal excitation. Using explicit SPH (smoothed-particle hydrodynamics), one can solve the governing equations for an incompressible fluid, simultaneously reproducing density fluctuations and describing the fluid in a many-particle system. We assume the following fluid-particle duality: a hydrodynamic interfacial tension between the inviscid and viscous components or a local interaction force between two types of fluid particles. The former can be induced in the horizontal direction when non-negligible non-uniformity of the particles occurs during forced 2D rotation, and the latter is non-negligible when the former is negligible. We performed a large-scale simulation of 3D liquid helium forced to rotate horizontally using 32 GPUs. Compared with the low-resolution calculation using 2.4 million particles, the high-resolution calculation using 19.6 million particles showed spinning vortices close to those of the theoretical solution. In the future, the introduction of quantum mechanical recombination among vortices will capture the behavior of quantum fluids. We obtained a promising venue to establish a practical simulation method for bulk liquid helium-4.
@@ Acknowledgment
Appendix
To facilitate understanding, we provide a schematic list of the equations for the derivation of the Navier–Stokes equations with spin angular momentum conservation in Fig. 9. We also provide a schematic summary of SPH in Fig. 10.
Figure 9: A schematic list of the equations for the derivation of the Navier–Stokes equations with spin angular momentum conservation in the literature [9]: (A) Cauchy’s equation of motion, (B) the conservation law of angular momentum, (C) the local expression of (B) using Reynolds’ transport theorem, (D) a decomposition of the angular momentum, (E) the evolution equation of $\bm{\mathbf{I}}$ obtained from (A), (C) and (D), (F) two components of the stress tensor, and (D), (F) two components of the stress tensor, (G) assumption of Newtonian fluid for the stress tensor and (H) for the coupled stress tensor, (I) contribution of the antisymmetric part of the stress tensor, and the resulting equations in (J), (K), and (L).
Figure 10: A schematic summary of SPH (smoothed-particle hydrodynamics): (A) the basic concept of SPH, (B) a list of SPH models used in this study, and (C) examples of kernel functions.